Dr. Bryan Carpenter

Department of Electronics and Computer Science University of Southampton Highfield Southampton S09 5NH

October 27, 1993

TAPELIB, Atlas Computer Laboratory, Rutherford Laboratory, Chilton Didcot OX11 OQX

Dear Tapelib,

the enclosed tape is

  USERID: cra
  ACCOUNT: cray16
  NAMES OF ENCLOSED EXABYTES: elc035

(This tape replaces the corrupted version of elc035, returned by you a few weeks ago).

Thank you.